# U-DiTs: Downsample Tokens in U-Shaped Diffusion Transformers

### v1

@@ Abstract
Diffusion Transformers (DiTs) introduce the transformer architecture to diffusion tasks for latent-space image generation. With an isotropic architecture that chains a series of transformer blocks, DiTs demonstrate competitive performance and good scalability; but meanwhile, the abandonment of U-Net by DiTs and their following improvements is worth rethinking. To this end, we conduct a simple toy experiment by comparing a U-Net architectured DiT with an isotropic one. It turns out that the U-Net architecture only gain a slight advantage amid the U-Net inductive bias, indicating potential redundancies within the U-Net-style DiT. Inspired by the discovery that U-Net backbone features are low-frequency-dominated, we perform token downsampling on the query-key-value tuple for self-attention and bring further improvements despite a considerable amount of reduction in computation. Based on self-attention with downsampled tokens, we propose a series of U-shaped DiTs (U-DiTs) in the paper and conduct extensive experiments to demonstrate the extraordinary performance of U-DiT models. The proposed U-DiT could outperform DiT-XL/2 with only 1/6 of its computation cost. Codes are available at https://github.com/YuchuanTian/U-DiT.
@@ 1 Introduction
Interestingly, DiTs have discarded the U-Net architecture [26] that is universally applied in manifold previous works, either in pixel [17; 11] or latent space [25]. The use of isotropic architectures in DiTs is indeed successful, as scaled-up DiT models achieve supreme performance. However, the abandonment of the widely-applied U-Net architecture by DiTs and their improvements [16; 8; 22] on latent-space image generation tasks triggers our curiosity, because the U-Net inductive bias is always believed to help denoising. Hence, we rethink deploying DiTs on a canonical U-Net architecture.
@@
Hence, we rethink the self-attention mechanism in DiT-UNet. The backbone in a latent U-Net denoiser provides a feature where low-frequency components dominate [27]. The discovery implies the existence of redundancies in backbone features: the attention module in the U-Net diffuser should highlight low-frequency domains. As previous theories praised downsampling for filtering high-frequency noises in diffusion [35], we seek to leverage this natural low-pass filter by performing token downsampling on the features for self-attention. Unlike previous transformer works [15; 38; 28] that downsample key-value pairs only, we radically downsample the query-key-value tuple altogether, such that self-attention is performed among downsampled latent tokens. It is surprising that when we incorporate self-attention with downsampled tokens into DiT-UNet, better results are achieved on latent U-Net diffusers with a significant reduction of computation.
@@ 2 Preliminaries
Vision Transformers. ViTs [13] have introduced a transformer backbone to vision tasks by patchifying the input and viewing an image as a sequence of patch tokens and have proved its effectiveness on large-scale image classification tasks. While ViTs adopt an isotropic architecture, some following works on vision transformers [33; 21] propose a pyramid-like hierarchical architecture that gradually downsamples the feature. The pyramid architecture has proved highly effective in classification and other downstream tasks.
Vision transformers are also mainstream backbones for denoising models. IPT [6] introduces an isotropic transformer backbone for denoising and other low-level tasks. Some later works [19; 18; 7] follow the isotropic convention, but other denoising works [34; 36] shift to U-Net backbones as their design. The pioneering work of U-ViT [1] and DiT [24] introduces full-transformer backbones to diffusion as denoisers.
Recent Advancements in Diffusion Transformers. Following DiTs, some works investigate the training and diffusion [14; 23] strategies of Diffusion Transformers. Other works focus on the design of the DiT backbone. DiffiT [16] introduces a new fusion method for conditions; FiT [22] and VisionLLaMA [8] strengthens DiT by introducing LLM tricks including RoPE2D [30] and SwishGLU. These transformer-based diffusion works agree on adopting isotropic architectures on latents, i.e. the latent feature space is not downsampled throughout the whole diffusion model. The authors of DiT [24] even regard the inductive bias of U-Net as “not crucial”.
U-Nets for Diffusion. From canonical works [17; 29; 11; 25], the design philosophy of U-Net [26] is generally accepted in diffusion. Specifically, Stable Diffusion [25] uses a U-Net-based denoiser on the compressed latent space for high-resolution image synthesis, which is highly successful in manifold generative tasks. Some previous trials on diffusion transformers [4; 16; 9] also adopt U-Net on pixel-space generation tasks; but strangely, they shifted to isotropic DiT-like structures for latent-space diffusion. Despite its popularity in pixel-space diffusion, the U-Net architecture is not widely accepted in recent transformer-oriented works on latent-space diffusion.
@@ 3 Investigating U-Net DiTs in Latent
+ Token Downsampling 0.90 89.43 21.36 15.13 0.29 0.44
@@
In fact, attention to downsampled tokens is not new. Previous works regarding vision transformers [15; 38] have proposed methods to downsample key-value pairs for computation cost reduction. Recent work on training-free acceleration of diffusion [28] also applies key-value downsampling on Stable Diffusion models. But these works maintain the number of queries, and thus the downsampling operation is not completely performed. Besides, these downsampling measures usually involves a reduction of tensor size, which could result in a significant loss in information.
@@ 4 Scaling the Model Up
Apart from the self-attention on downsampling as introduced in the toy experiment (Section 3), we further introduce a series of modifications to U-DiTs, including cosine similarity attention [20; 18], RoPE2D [30; 22; 8], depthwise conv FFN [34; 3; 38], and re-parametrization [12; 31]. The contribution of each modification is quantitatively evaluated in Sec. 6.
@@ 4.1 U-DiT at Larger Scales
Comparison with DiTs and their improvements. In order to validate the effectiveness of the proposed U-DiT models beyond simple toy experiments, we scale them up and compare them with DiTs [24] of larger sizes. For a fair comparison, we use the same sets of training hyperparameters as DiT; all models are trained for 400K iterations. The results on ImageNet 256$\times$256 are shown in Tab. 2, where we scale U-DiTs to $\sim 6e9$, $\sim 20e9$, $\sim 80e9$ FLOPs respectively and compare them with DiTs of similar computation costs.
@@ 4.2 Ablations
Next, we inspect other modifications that further refine U-DiTs (lower part of Tab. 6). Swin Transformer V2 [20] proposes a stronger variant of self-attention: instead of directly multiplying Q and K matrices, cosine similarities between queries and keys are used. We apply the design to our self-attention, which yields $\sim 2.5$FIDs of improvement. RoPE [30] is a powerful positional embedding method, which has been widely applied in Large Language Models. Following the latest diffusion transformer works [22; 8], we inject 2-dimensional RoPE (RoPE2D) into queries and keys right before self-attention. The introduction of RoPE2D improves performance by $\sim 2.5$FIDs. Some recent transformer works strengthen MLP by inserting a depthwise convolution layer between two linear mappings [34; 3; 38]. As the measure is proved effective in these works, we borrow it to our U-DiT model, improving $\sim 5$FIDs. As re-parametrization during training [12] could improve model performance, we apply the trick to FFN [31] and bring an additional improvement of $\sim 3.5$FIDs.
Above all, based on the components mentioned above, the proposed U-DiTs could outcompete plain DiT-UNets and isotropic DiTs by large margins.

### v2

@@ Abstract
Diffusion Transformers (DiTs) introduce the transformer architecture to diffusion tasks for latent-space image generation. With an isotropic architecture that chains a series of transformer blocks, DiTs demonstrate competitive performance and good scalability; but meanwhile, the abandonment of U-Net by DiTs and their following improvements is worth rethinking. To this end, we conduct a simple toy experiment by comparing a U-Net architectured DiT with an isotropic one. It turns out that the U-Net architecture only gain a slight advantage amid the U-Net inductive bias, indicating potential redundancies within the U-Net-style DiT. Inspired by the discovery that U-Net backbone features are low-frequency-dominated, we perform token downsampling on the query-key-value tuple for self-attention that bring further improvements despite a considerable amount of reduction in computation. Based on self-attention with downsampled tokens, we propose a series of U-shaped DiTs (U-DiTs) in the paper and conduct extensive experiments to demonstrate the extraordinary performance of U-DiT models. The proposed U-DiT could outperform DiT-XL/2 with only 1/6 of its computation cost. Codes are available at https://github.com/YuchuanTian/U-DiT.
@@ 1 Introduction
Interestingly, DiTs have discarded the U-Net architecture [27] that is universally applied in manifold previous works, either in pixel [18; 12] or latent space [26]. The use of isotropic architectures in DiTs is indeed successful, as scaled-up DiT models achieve supreme performance. However, the abandonment of the widely-applied U-Net architecture by DiTs and their improvements [17; 9; 23] on latent-space image generation tasks triggers our curiosity, because the U-Net inductive bias is always believed to help denoising. Hence, we rethink deploying DiTs on a canonical U-Net architecture.
@@
Hence, we rethink the self-attention mechanism in DiT-UNet. The backbone in a latent U-Net denoiser provides a feature where low-frequency components dominate [28]. The discovery implies the existence of redundancies in backbone features: the attention module in the U-Net diffuser should highlight low-frequency domains. As previous theories praised downsampling for filtering high-frequency noises in diffusion [36], we seek to leverage this natural low-pass filter by performing token downsampling on the features for self-attention. Unlike previous transformer works [16; 39; 29] that downsample key-value pairs only, we radically downsample the query-key-value tuple altogether, such that self-attention is performed among downsampled latent tokens. It is surprising that when we incorporate self-attention with downsampled tokens into DiT-UNet, better results are achieved on latent U-Net diffusers with a significant reduction of computation.
@@ 2 Preliminaries
Vision Transformers. ViTs [14] have introduced a transformer backbone to vision tasks by patchifying the input and viewing an image as a sequence of patch tokens and have proved its effectiveness on large-scale image classification tasks. While ViTs adopt an isotropic architecture, some following works on vision transformers [34; 22] propose a pyramid-like hierarchical architecture that gradually downsamples the feature. The pyramid architecture has proved highly effective in classification and other downstream tasks.
Vision transformers are also mainstream backbones for denoising models. IPT [6] introduces an isotropic transformer backbone for denoising and other low-level tasks. Some later works [20; 19; 8] follow the isotropic convention, but other denoising works [35; 37] shift to U-Net backbones as their design. The pioneering work of U-ViT [1] and DiT [25] introduces full-transformer backbones to diffusion as denoisers.
Recent Advancements in Diffusion Transformers. Following DiTs, some works investigate the training and diffusion [15; 24] strategies of Diffusion Transformers. Other works focus on the design of the DiT backbone. DiffiT [7; 17] introduces a new fusion method for conditions; FiT [23] and VisionLLaMA [9] strengthens DiT by introducing LLM tricks including RoPE2D [31] and SwishGLU. These transformer-based diffusion works agree on adopting isotropic architectures on latents, i.e. the latent feature space is not downsampled throughout the whole diffusion model. The authors of DiT [25] even regard the inductive bias of U-Net as “not crucial”.
U-Nets for Diffusion. From canonical works [18; 30; 12; 26], the design philosophy of U-Net [27] is generally accepted in diffusion. Specifically, Stable Diffusion [26] uses a U-Net-based denoiser on the compressed latent space for high-resolution image synthesis, which is highly successful in manifold generative tasks. Some previous trials on diffusion transformers [4; 17; 10] also adopt U-Net on pixel-space generation tasks; but strangely, they shifted to isotropic DiT-like structures for latent-space diffusion. Despite its popularity in pixel-space diffusion, the U-Net architecture is not widely accepted in recent transformer-oriented works on latent-space diffusion.
@@ 3 Investigating U-Net DiTs in Latent
DiT-UNet+Key-Value Downsampling 0.91 94.38 23.21 14.32 0.27 0.40
DiT-UNet+Token Downsampling (Ours) 0.90 89.43 21.36 15.13 0.29 0.44
@@
In fact, attention to downsampled tokens is not new. Previous works regarding vision transformers [16; 39] have proposed methods to downsample key-value pairs for computation cost reduction. Recent work on training-free acceleration of diffusion [29] also applies key-value downsampling on Stable Diffusion models. But these works maintain the number of queries, and thus the downsampling operation is not completely performed. Besides, these downsampling measures usually involves a reduction of tensor size, which could result in a significant loss in information.
@@ 4 Scaling the Model Up
Apart from the self-attention on downsampling as introduced in the toy experiment (Section 3), we further introduce a series of modifications to U-DiTs, including cosine similarity attention [21; 19], RoPE2D [31; 23; 9], depthwise conv FFN [35; 3; 39], and re-parametrization [13; 32]. The contribution of each modification is quantitatively evaluated in Sec. 8.
@@ 4.1 U-DiT at Larger Scales
Comparison with DiTs and their improvements. In order to validate the effectiveness of the proposed U-DiT models beyond simple toy experiments, we scale them up and compare them with DiTs [25] of larger sizes. For a fair comparison, we use the same sets of training hyperparameters as DiT; all models are trained for 400K iterations. The results on ImageNet 256$\times$256 are shown in Tab. 2, where we scale U-DiTs to $\sim 6e9$, $\sim 20e9$, $\sim 80e9$ FLOPs respectively and compare them with DiTs of similar computation costs, more details about the U-DiT architectures are shown in Tab. 7.
@@
Larger image size. We additionally compare the generation performance of U-DiT-B and DiT-XL/2 on ImageNet 512 $\times$ 512 under exactly the same training setting. As shown in Tab. 5, U-DiT-B could still outcompete DiT-XL/2 that is approximately 5 times larger in FLOPs.
@@
ImageNet 512$\times$512
[…]
DiT-XL/2 524.7 20.94 6.78 66.30 0.745 0.581
U-DiT-B 106.7 15.39 6.86 92.73 0.756 0.605
Table 5: Comparing U-DiTs against DiTs on ImageNet 512$\times$512 generation. Experiments with a supermark are replicated according to the official code of DiT. We compare models trained for 400K iterations with the standard training hyperparameters of DiT.
@@ 4.2 Ablations
Next, we inspect other modifications that further refine U-DiTs (lower part of Tab. 8). Swin Transformer V2 [21] proposes a stronger variant of self-attention: instead of directly multiplying Q and K matrices, cosine similarities between queries and keys are used. We apply the design to our self-attention, which yields $\sim 2.5$FIDs of improvement. RoPE [31] is a powerful positional embedding method, which has been widely applied in Large Language Models. Following the latest diffusion transformer works [23; 9], we inject 2-dimensional RoPE (RoPE2D) into queries and keys right before self-attention. The introduction of RoPE2D improves performance by $\sim 2.5$FIDs. Some recent transformer works strengthen MLP by inserting a depthwise convolution layer between two linear mappings [35; 3; 39]. As the measure is proved effective in these works, we borrow it to our U-DiT model, improving $\sim 5$FIDs. As re-parametrization during training [13] could improve model performance, we apply the trick to FFN [32] and bring an additional improvement of $\sim 3.5$FIDs. Above all, based on these components, the proposed U-DiTs are further improved.
Apart from the modifications that improve U-DiT, it is worth noting that vanilla U-DiTs (i.e. U-DiTs without any of the modifications mentioned above) are still competitive. According to Tab. 9, vanilla U-DiT-L could still achieve $\sim 8$FIDs of advantage over DiT-XL/2.
Model Params (M) FLOPs (G) Channel Head Number Encoder-Decoder
U-DiT-S 52.05 6.04 96 4 [2, 5, 8, 5, 2]
U-DiT-B 204.43 22.22 192 8 [2, 5, 8, 5, 2]
U-DiT-L 810.19 85.00 384 16 [2, 5, 8, 5, 2]
Table 7: Configurations of U-DiTs architecture with different model sizes. Channel represents the initial output channel number of first layer. Encoder-Decoder denotes the transformer block number of encoder and decoder module.
@@
ImageNet 256$\times$256
[…]
U-DiT-S (Vanilla) 5.91 41.01 10.96 39.29 0.489 0.622
U-DiT-S (+All Mods) 6.04 31.51 8.97 51.62 0.543 0.633
U-DiT-B (Vanilla) 21.96 20.89 7.33 72.85 0.611 0.637
U-DiT-B (+All Mods) 22.22 16.64 6.33 85.15 0.642 0.639
U-DiT-L (Vanilla) 84.48 12.04 5.37 102.63 0.684 0.628
U-DiT-L (+All Mods) 85.00 10.08 5.21 112.44 0.702 0.631
Table 9: Comparison between vanilla U-DiTs and improved U-DiTs with all modifications. With negligible extra computational overhead, the proposed modifications could improve the performance of U-DiT; but it is worth noting that vanilla U-DiTs are powerful enough against DiTs.
@@ 5 Conclusion
Broader Impacts. Due to the biases in the training data set, the generated content may contain pornographic, racist, hate and violent information. But we emphasize that the potential for misuse is mitigated through vigilant application.
Discussion of Safeguards. For cautionous usage, we suggest an algorithm capable of checking generated images, in order to identify and mitigate content that contravenes legal or ethical usages.